1. Express the rate of reaction (in differential form) with respect to each species in the following reactions (reactants and products): (3 points)

a. N2 + 3 H2 → 2 NH3

Rate =

b. C3H8 + 5 O2 → 3 CO2 + 4 H2O

Rate = c.

3 Ca(NO3)2 + 2 Li3PO4 → 6 LiNO3 + Ca3(PO4)2

Rate =
Answer #1

a. N2 + 3 H2 → 2 NH3

dN2/dt = 1/3*d(H2)/dt

-dN2/dt = 1/2*d(NH3)/dt

-d(H2)/dt = 2/3*d(NH3)/dt

b. C3H8 + 5 O2 → 3 CO2 + 4 H2O

d(C3H8)/dT = 1/5*​d(O2)/dT

-d(C3H8)/dT = 1/3*​d(CO2)/dT

-d(C3H8)/dT = 1/4*​d(H2O)/dT

c)

3 Ca(NO3)2 + 2 Li3PO4 → 6 LiNO3 + Ca3(PO4)2

d(Ca(NO3)2)/dT = 3/2* d(Li3PO4)/dt

-d(Ca(NO3)2)/dT = 3/6* d(LiNO3 )/dt

-d(Ca(NO3)2)/dT = 3*d(Ca3(PO4)2)/dt